# Analysis and Simulation of a Fluid-Heat System in a Thin, Rough Layer in Contact With a Solid Bulk Domain

### v2

@@ 2.2 Underlying model equations
[…]
@@
[…]
where $\llbracket\theta\rrbracket=(\theta^{f}-\theta^{s})$ denotes the usual jump across an interface, $\alpha$ denotes the heat exchange parameter, and $f_{\varepsilon}$ describes a heat source. Here, the chosen heat source is motivated by our underlying application problem of grinding, there the system heats up through friction between the workpiece and the grinding wheel. Therefore, the heat is produced at the interface. To simplify the notation the heat source is only located on the interface side of the solid domain. Through the Robin condition heat is then also transferred into the fluid.
Note, the specific $\varepsilon$-scaling can be reasoned from a non-dimensionalization step and can be found in similar problems [20, 29] and originates from the scaling of the physical parameters (e.g., a fast heat conductivity in the thin layer compared to the height of the layer and the scaling of the density keeps the total amount of heat energy constant with respect to $\varepsilon$). Clearly Eq. 2.1b could be multiplied with $\varepsilon$ and the scaling would then only enter on the right-hand side of (2.1f). For the following analysis, the above form is more appropriate.
@@
[…]
@@ 2.3 Assumptions on data
    For the heat source it holds $f_{\varepsilon}\in L^{\infty}(S\times\Gamma_{\varepsilon})$ and $C_{f}\coloneqq\sup_{\varepsilon>0}\|f_{\varepsilon}\|_{L^{\infty}(S\times\Gamma_{\varepsilon})}<\infty$.
@@
    For the prescribed velocity on the lateral fluid boundaries, we assume there is $u_{BC}\in W^{1,\infty}(\partial\Sigma\times(0,1))^{d}$ with the properties:
@@
    • $u_{BC}(\tilde{x},0)=u_{\text{motion}}$, where it still holds that $u_{\text{motion}}\cdot e_{d}=0$,
@@
    • inflow and outflow conditions $\left(u_{BC}\cdot n_{\Sigma}\right)_{|\Sigma^{\text{in}}}<0\text{ and }\left(u_{BC}\cdot n_{\Sigma}\right)_{|\Sigma^{\text{out}}}\geq 0$.
@@
    $\int_{\partial\Sigma\times(0,1)}u_{BC}\cdot n_{\Sigma}\,\mathrm{d}\sigma_{x}=0$ (2.2)
    to ensure the existence of solutions. Here, $n_{\Sigma}$ is the normal vector that points outward from $\Sigma\times(0,1)$. The boundary condition on the thin layer is given by $u_{\varepsilon,BC}(x)=u_{BC}(\tilde{x},x_{d}/\varepsilon)$ for $x\in\Sigma^{\text{in}}_{\varepsilon}\cup\Sigma^{\text{out}}_{\varepsilon}$.
@@
    For the initial conditions there are limit functions $\theta^{s}_{0}\in L^{2}(\Omega)$ and $\theta^{f}_{0}\in L^{2}(\Sigma)$ such that $\chi_{\Omega^{s}_{\varepsilon}}\theta_{\varepsilon,0}\to\theta^{s}_{0}$ in $L^{2}(\Omega^{s}_{\varepsilon})$ and $\chi_{\Omega^{f}_{\varepsilon}}\theta^{f}_{\varepsilon,0}\stackrel{{\scriptstyle 2}}{{\rightharpoonup}}\chi_{Z}\theta^{f}_{0}$ in $L^{2}(\Omega_{\varepsilon})$, for $\varepsilon\to 0$. Here, $\stackrel{{\scriptstyle 2}}{{\rightharpoonup}}$ denotes two-scale convergence in the thin layer which will be introduced in Definition 1.
@@
    For the heat source there is a $f\in L^{2}(S\times\Sigma\times\Gamma)$ such that $f_{\varepsilon}\stackrel{{\scriptstyle 2}}{{\rightharpoonup}}f$, for $\varepsilon\to 0$.
@@
Given the assumption (A5), we can always construct a continuation $U\in H^{1}(\Sigma\times(0,1))$ of $u_{BC}$, with $U(\tilde{x},y)=0$ for $y\geq\gamma_{0}$ and $\operatorname{div}{U}=0$, [32, Theorem. 3.5]. By setting $U_{\varepsilon}(x)=(U_{1},\dots,\varepsilon U_{d})(\tilde{x},\frac{x_{d}}{\varepsilon})$ the extension can also be transferred to the thin rough layer $\Omega^{f}_{\varepsilon}$, this construction was already derived in [12]. We expect that the condition $u_{BC}=0$ for $x_{d}\geq\gamma_{0}$ could be removed and the extension still be carried out by first building an extension around the boundary falling back to a rectangle of height $\varepsilon\gamma_{0}$ and then proceeding as above, but this is not further investigated here. Examples of functions that fulfill the assumptions can be found in Section 5. We could also utilize a time-dependent boundary function $u_{BC}\in L^{\infty}(S;W^{1,\infty}(\partial\Sigma\times(0,1)))^{d}$, this would not change the following results. To simplify the notation we keep a time-independent $u_{BC}$. Additionally, given the prescribed boundary conditions of $u_{BC}$, it holds $u_{\varepsilon,BC}=u_{\text{motion}}$ on $\Sigma$ and we and we have inflow at $\Sigma^{\text{in}}_{\varepsilon}$ and outflow at $\Sigma^{\text{out}}_{\varepsilon}$, to be precise
$\left(u_{\varepsilon,BC}\cdot n_{\varepsilon}\right)_{|\Sigma^{\text{in}}_{\varepsilon}}<0\quad\text{ and }\quad\left(u_{\varepsilon,BC}\cdot n_{\varepsilon}\right)_{|\Sigma^{\text{out}}_{\varepsilon}}\geq 0.$ (2.3)
@@
We provide some additional comments regarding our model and our mathematical assumptions specifically in regards to the application case of grinding:
@@
    During the actual physical process there is, of course, contact between the grinding wheel and the workpiece. In our model, that would imply that $\Omega^{s}_{\varepsilon}$ and $\Sigma$ are touching (i.e., $\gamma_{0}=0$). Note that for dimension $d=2$, the model would no longer be reasonable as any contact would necessarily disrupt fluid flow; while for $d=3$ it would still be possible for the fluid to flow around the contact points. In other works where the behavior of the fluid pressure inside the grinding gap is studied, an artificial gap is introduced such that the pressure can be modeled by the Reynolds equation [53, 55]. Their findings demonstrate that this is a reasonable approximation to the real dynamics. Therefore, we keep this simplification of no direct contact and focus on how far a temperature-dependent viscosity influences the model.
@@
    The heat which is released during grinding is caused by the friction between the grains of the grinding wheel and the workpiece and only part of this heat is released into the system directly (e.g., part of it is also transported with the formation of small chips [48]). Here, we consider a simplified model where we assume a prescribed heat source at the interface but where we do not account for any mechanical effects in the solid part (in particular: no changing geometries, no chips, and no strains).
@@
    The relationship between temperature and viscosity of fluids is complex and often modeled heuristically. A typical example is the Vogel-Fulcher-Tammann equation given as
@@
    where $\mu_{0},\alpha,T_{0}$ are fitting parameters, we refer to [30] and references therein. This, of course, is only valid in a certain temperature range (usually a proper subset of the liquid phase range of the substance). However, it suggests that the relationship tends to be very smooth and that the Lipschitz continuity is a reasonable assumption.
  • $(iv)$
    The specific assumptions regarding $u_{BC}$ and, as a consequence, $u_{\varepsilon,BC}$ are a bit technical, but they ensure that there is an inflow and an outflow region and that $\Sigma$ and $\Gamma_{\varepsilon}$ are in relative motion to each other. The regularity assumption ensures compatibility across the different boundary parts. The inflow condition models the continuous injection of cooling liquid where the corresponding Dirichlet condition Eq. 2.1d accounts for the cooling effect.
  • $(v)$
    We do not include inertia, turbulence, and buoyancy effects by using the quasi-stationary Stokes system. This is somewhat justified by the thinness of the contact area where the fluid flow does not have enough space to create strong turbulence and is mainly pulled along by the moving wheel as demonstrated in the experiments in [21]. We also neglect the real and important phenomenon of air entrainment into the fluid layer [34, 36]. Additionally, the fluid movement adjusts on a shorter time interval than the temperature.
3 The microscale model
@@ 3.1 Auxiliary results
    For all $\varepsilon>0$, there exist a linear extension operator $\mathbb{E}_{\varepsilon}:W^{1,p}(\Omega^{f}_{\varepsilon})\to W^{1,p}(\Omega_{\varepsilon})$ such that
    $\|\mathbb{E}_{\varepsilon}\psi\|_{W^{1,p}(\Omega_{\varepsilon})}\leq C\|\psi\|_{W^{1,p}(\Omega^{f}_{\varepsilon})}\quad\text{ for all }\psi\in W^{1,p}(\Omega^{f}_{\varepsilon})$
    and for all $p\in[1,\infty)$. Here, $C>0$ may depend on $p$ but is independent of $\varepsilon$.
@@
    There exists a diffeomorphism $\Phi:\Omega\to\Omega^{s}_{\varepsilon}$ such that
    $\frac{1}{C}\|\varphi\|_{W^{1,p}(\Omega^{s}_{\varepsilon})}\leq\|\varphi\circ\Phi\|_{W^{1,p}(\Omega)}\leq C\|\varphi\|_{W^{1,p}(\Omega^{s}_{\varepsilon})}\quad\text{ for all }\varphi\in W^{1,p}(\Omega^{s}_{\varepsilon})$
    and for all $p\in[1,\infty)$. Here, $C>0$ may depend on $p$ but is independent of $\varepsilon$.
@@
$(i)$. The extension operators can be constructed using a classical extension procedure [1]. For thin layers with a more general geometry, this was carried out in detail and further studied in [28, Theorem 1]. The general idea is to first construct extension operators for a single cell, in our case from $W^{1,p}(Z)$ to $W^{1,p}(Y^{d})$. Those exist since $\gamma$ is Lipschitz-continuous, and they are naturally $\varepsilon$-independent (as there is no $\varepsilon$ involved yet). The statement itself then follows by gluing these extension operators together via a periodic partition of unity in conjunction with a scaling argument.
$(ii)$. For the transformation of $\Phi$ to the complete domain $\Omega$, we define the functions
$\phi_{1}(x)\coloneqq\varepsilon\gamma(\nicefrac{{\tilde{x}}}{{\varepsilon}})+x_{d}\frac{H-\varepsilon\gamma(\nicefrac{{\tilde{x}}}{{\varepsilon}})}{H}\quad\text{ and }\quad\phi_{2}(x)\coloneqq H\frac{x_{d}-\varepsilon\gamma(\nicefrac{{\tilde{x}}}{{\varepsilon}})}{H-\varepsilon\gamma(\nicefrac{{\tilde{x}}}{{\varepsilon}})},$
where for a fixed $\tilde{x}\in\Sigma$ it holds
$\phi_{1}(\tilde{x},\cdot):(0,H)\to(\varepsilon\gamma(\nicefrac{{\tilde{x}}}{{\varepsilon}}),H)\quad\text{ and }\quad\phi_{2}(\tilde{x},\cdot):(\varepsilon\gamma(\nicefrac{{\tilde{x}}}{{\varepsilon}}),H)\to(0,H).$
With $\phi_{1}$ we obtain the transformation
$\Phi:\Omega\to\Omega^{s}_{\varepsilon},\quad x\mapsto(\tilde{x},\phi_{1}(\tilde{x},x_{d}))$
Since $\gamma$ is in $C^{1}(\overline{Y^{d-1}};[\gamma_{0},1])$ and $H>\varepsilon$ it holds that $G$ is bijective and continuously differentiable. The inverse is given by $\Phi^{-1}(x)=(\tilde{x},\phi_{2}(\tilde{x},x_{d}))$. Given the properties of $\gamma$, it holds
$|\phi_{1}(x)|,|\phi_{2}(y)|\leq C,\quad|\nabla\phi_{1}(x)|,|\nabla\phi_{2}(y)|\leq C\quad\text{ and }\quad c\leq\det(\nabla\Phi),\det(\nabla\Phi^{-1})\leq C$
for all $x\in\Omega,y\in\Omega^{s}_{\varepsilon}$, where $C,c>0$ can be chosen independent of $\varepsilon$. From these bounds we also obtain $\varepsilon$ independent bounds of $|\nabla\Phi|$ and $|\nabla\Phi^{-1}|$. The estimates in the $W^{1,p}$ norm for $\varphi$ and $\varphi\circ\Phi$ then follow via integration by substitution. ∎
@@
    $\|\psi\|_{L^{p}(\Gamma_{\varepsilon})}\leq\frac{C}{\varepsilon^{1/p}}\|\psi\|_{W^{1,p}(\Omega^{f}_{\varepsilon})}\quad\text{for all}\quad\psi\in W^{1,p}(\Omega^{f}_{\varepsilon})$
@@
    $\|\phi\|_{L^{p}(\Gamma_{\varepsilon})}\leq C\|\phi\|_{W^{1,p}(\Omega^{s}_{\varepsilon})}\quad\text{for all}\quad\varphi\in W^{1,p}(\Omega^{s}_{\varepsilon})$
@@
$(i)$. This estimate for $\psi\in W^{1,p}(\Omega^{f}_{\varepsilon})$ is quite standard and follows by a splitting argument, which we shortly sketch. Split the integral over $\Gamma_{\varepsilon}$ into integrals over the boundary of the cell $\varepsilon Z$, applying a scaling argument followed by a trace estimate in the rescaled domain and then transforming back to the $\varepsilon$ domain. Note, that the estimate could be improved by a $\varepsilon^{1/p}$ in front of the gradient terms, but this is not needed in the following analysis.
$(ii)$. A similar estimate was derived in [19, Proposition 2], albeit with homogeneous boundary conditions at the outer boundary $\partial\Omega^{s}_{\varepsilon}\setminus\Gamma_{\varepsilon}$ for $\varphi\in H^{1}(\Omega^{s}_{\varepsilon})$. By construction of the transformation $\Phi:\Omega\to\Omega^{s}_{\varepsilon}$ from Lemma 1, it holds $(\varphi\circ\Phi)_{|\Sigma}=\varphi_{|\Gamma_{\varepsilon}}$ and
$\int_{\Gamma_{\varepsilon}}|\varphi(x)|^{p}\,\mathrm{d}\sigma_{x}$ $=\int_{\Sigma}\left|\varphi(\tilde{x},\varepsilon\gamma(\nicefrac{{\tilde{x}}}{{\varepsilon}}))\right|^{p}\sqrt{1+|\varepsilon\nabla_{\tilde{x}}\left[\gamma(\nicefrac{{\tilde{x}}}{{\varepsilon}})\right]|^{2}}\,\mathrm{d}\sigma_{\tilde{x}}$
$=\int_{\Sigma}\left|(\varphi\circ\Phi)(\tilde{x},0)\right|^{p}\sqrt{1+|\left[\nabla_{\tilde{x}}\gamma\right](\nicefrac{{\tilde{x}}}{{\varepsilon}})|^{2}}\,\mathrm{d}\sigma_{\tilde{x}}$
$\leq C\int_{\Sigma}\left|(\varphi\circ\Phi)(\tilde{x},0)\right|^{p}\,\mathrm{d}\sigma_{\tilde{x}}.$
On $\Omega$ we can use a trace theorem independent of $\varepsilon$ to obtain with Lemma 1.($ii$)
$\|\varphi\|_{L^{p}(\Gamma_{\varepsilon})}\leq C\|\varphi\circ\Phi\|_{L^{p}(\Sigma)}\leq C\|\varphi\circ\Phi\|_{W^{1,p}(\Omega)}\leq C\|\varphi\|_{W^{1,p}(\Omega^{s}_{\varepsilon})}.$
$(iii).$ This follows by a scaling argument. With the extension operator from Lemma 1, it holds
$\|\psi\|_{L^{p}(\Sigma^{f}_{\varepsilon})}$ $=\|\mathbb{E}\psi\|_{L^{p}(\Sigma^{f}_{\varepsilon})}\leq\|\mathbb{E}\psi\|_{L^{p}(\partial\Sigma\times(0,\varepsilon))}=\varepsilon^{\frac{1}{p}}\|\mathbb{E}\psi(\tilde{x},\varepsilon x_{d})\|_{L^{p}(\partial\Sigma\times(0,1))}$
$\leq C\varepsilon^{\frac{1}{p}}\|\mathbb{E}\psi(\tilde{x},\varepsilon x_{d})\|_{W^{1,p}(\Omega)}\leq C\|\mathbb{E}\psi\|_{W^{1,p}(\Omega_{\varepsilon})}\leq C\|\psi\|_{W^{1,p}(\Omega^{f}_{\varepsilon})}.$
Where we used the trace theorem on $\Omega$. Note, as before, the estimate could be improved by a $\varepsilon^{1/p}$ in front of the term with the $\partial_{x_{d}}$ norm, but this is not needed in the following analysis. ∎
@@
For any $\psi\in L^{\infty}(S;L^{2}(\Omega_{\varepsilon})\cap L^{2}(S;H^{1}(\Omega_{\varepsilon}))$, let $\bar{\psi}\in L^{\infty}(S;L^{2}(\Omega))\cap L^{2}(S;H^{1}(\Omega))$ be the corresponding upscaled function given by $\bar{\psi}(t,\tilde{x},x_{d})=\psi(t,\tilde{x},\varepsilon x_{d})$. Using standard interpolation embeddings (see, e.g., [41, Chapter 2, Inequality 3.8]), it holds

### v3

@@ 2.2 The micro model
$\theta_{\varepsilon}$ $=0$ $\text{ on }S\times\Sigma^{\text{in}}_{\varepsilon},$ (1d)
$-\kappa\nabla\theta_{\varepsilon}\cdot n$ $=0$ $\text{ on }S\times\left(\Sigma^{\text{out}}_{\varepsilon}\cup\Sigma^{s}_{\varepsilon}\cup\Sigma\right).$ (1e)
@@
$-\frac{1}{\varepsilon}\kappa^{f}\nabla\theta^{f}_{\varepsilon}\cdot n^{f}$ $=\alpha\llbracket\theta_{\varepsilon}\rrbracket+f_{\varepsilon}^{f}\quad$ $\text{ on }S\times\Gamma_{\varepsilon},$ (1f)
$-\kappa^{s}\nabla\theta^{s}_{\varepsilon}\cdot n^{s}$ $=\alpha\llbracket\theta_{\varepsilon}\rrbracket+f_{\varepsilon}^{s}\quad$ $\text{ on }S\times\Gamma_{\varepsilon},$ (1g)
where $\llbracket\theta\rrbracket=(\theta^{f}-\theta^{s})$ denotes the usual jump across an interface, $\alpha$ denotes the heat exchange parameter, and $f_{\varepsilon}^{f,s}$ describe heat sources.
Note, the chosen $\varepsilon$-scaling appears in similar problems [41, 22] and is designed to keep the contributions of the equations in the effect model for $\varepsilon\to 0$. From a physical point of view, the scaling originates from assumptions about the physical parameters (e.g., a fast heat conductivity in the thin layer compared to the height of the layer) and can be incorporated by a non-dimensionalization step; see for example [13, Section 2.1], where this is explained for fluid flow through thin cracks.
@@
$-\operatorname{div}\left(\mu\left(\theta^{f}_{\varepsilon}\right)\nabla u_{\varepsilon}\right)+\nabla p_{\varepsilon}$ $=0\quad\quad$ $\text{ in }S\times\Omega^{f}_{\varepsilon},$ (1h)
$\operatorname{div}\left(u_{\varepsilon}\right)$ $=0\quad\quad$ $\text{ in }S\times\Omega^{f}_{\varepsilon},$ (1i)
$u_{\varepsilon}$ $=u_{\varepsilon,BC}$ $\text{ on }S\times\partial\Omega^{f}_{\varepsilon},$ (1j)
$\int_{\Omega^{f}_{\varepsilon}}p_{\varepsilon}\,\mathrm{d}x$ $=0$ $\text{ in }S.$ (1k)
Here $u_{\varepsilon,BC}\in H^{1/2}(\partial\Omega^{f}_{\varepsilon})\cap L^{\infty}(\partial\Omega^{f}_{\varepsilon})$ describes a given boundary flow.
@@ 2.3 Assumptions on data
    For the heat sources it holds $f_{\varepsilon}^{f,s}\in L^{\infty}(S\times\Gamma_{\varepsilon})$ and $\sup_{\varepsilon>0}\|f_{\varepsilon}^{f,s}\|_{L^{\infty}(S\times\Gamma_{\varepsilon})}<\infty$.
@@
    For the Dirichlet condition on the lateral fluid boundaries, we assume there is $u_{BC}\in H^{1/2}(\partial\Sigma\times(0,1))^{d}\cap L^{\infty}(\partial\Sigma\times(0,1))^{d}$ with the properties:
@@
    • $u_{BC}(\tilde{x},0)=u_{\text{motion}}$, with $u_{\text{motion}}\in\mathbb{R}^{d}$ and $u_{\text{motion}}\cdot e_{d}=0$,
@@
    • inflow and outflow conditions $\left(u_{BC}\cdot n_{\Sigma}\right)_{|\Sigma^{\text{in}}\times(0,1)}<0\text{ and }\left(u_{BC}\cdot n_{\Sigma}\right)_{|\Sigma^{\text{out}}\times(0,1)}\geq 0$.
@@
    $\int_{\partial\Sigma\times(0,1)}u_{BC}\cdot n_{\Sigma}\,\mathrm{d}\sigma_{x}=0$ (2)
    to ensure the existence of solutions. Here, $n_{\Sigma}$ is the normal vector that points outward from $\Sigma\times(0,1)$.
@@
    For the initial conditions there are limit functions $\theta^{s}_{0}\in L^{2}(\Omega)$ and $\theta^{f}_{0}\in L^{2}(\Sigma)$ such that $\chi_{\Omega^{s}_{\varepsilon}}\theta_{\varepsilon,0}\to\theta^{s}_{0}$ in $L^{2}(\Omega^{s}_{\varepsilon})$ and $\chi_{\Omega^{f}_{\varepsilon}}\theta^{f}_{\varepsilon,0}\stackrel{{\scriptstyle 2}}{{\rightharpoonup}}\chi_{Z}\theta^{f}_{0}$ in $L^{2}(\Omega_{\varepsilon})$, for $\varepsilon\to 0$. Here, $\stackrel{{\scriptstyle 2}}{{\rightharpoonup}}$ denotes two-scale convergence in the thin layer which is introduced in Definition 1.
@@
    For the heat sources there are $f^{f,s}\in L^{2}(S\times\Sigma\times\Gamma)$ such that $f_{\varepsilon}^{f,s}\stackrel{{\scriptstyle 2}}{{\rightharpoonup}}f^{f,s}$, for $\varepsilon\to 0$.
@@
The assumption (A5) characterizes the boundary condition on the scaled boundary $\partial\Sigma\times(0,1)$, the microscopic boundary condition is defined via
$u_{\varepsilon,BC}(x)\coloneqq\begin{cases}u_{BC}(\tilde{x},x_{d}/\varepsilon)\quad&\text{for }x\in\Sigma^{\text{in}}_{\varepsilon}\cup\Sigma^{\text{out}}_{\varepsilon},\\ u_{\text{motion}}&\text{for }x\in\Sigma,\\ 0&\text{for }x\in\Gamma_{\varepsilon}.\end{cases}$
The inflow and outlfow conditions of $u_{BC}$ are transferred to $u_{\varepsilon,BC}$, to be precise
$\left(u_{\varepsilon,BC}\cdot n\right)_{|\Sigma^{\text{in}}_{\varepsilon}}<0\quad\text{ and }\quad\left(u_{\varepsilon,BC}\cdot n\right)_{|\Sigma^{\text{out}}_{\varepsilon}}\geq 0.$ (3)
Examples of functions that fulfill the assumptions can be found in Section 5. Furthermore, by assumption (A5), we can construct a continuation $U\in H^{1}(\Sigma\times(0,1))$ of $u_{BC}$, with $U(\tilde{x},y)=0$ for $y\geq\gamma_{0}$ and $\operatorname{div}{U}=0$, [42, Theorem. 3.5]. By setting $U_{\varepsilon}(x)=(U_{1},\dots,\varepsilon U_{d})(\tilde{x},\frac{x_{d}}{\varepsilon})$ the extension can also be transferred to the thin rough layer $\Omega^{f}_{\varepsilon}$ (this requires $u_{BC}\cdot e_{d}=0$ such that $U_{\varepsilon}$ has the correct boundary condition), this construction was already derived in [25].
@@
We provide some additional comments regarding our model and our mathematical assumptions:
@@
    $u_{\text{motion}}$ describes a prescribed velocity at the bottom boundary (for example, the movement induced by a rotating grinding wheel in our motivation). It is not mandatory that $u_{\text{motion}}$ is constant but could also be function in $H^{1/2}(\Sigma)^{d}\cap L^{\infty}(\Sigma)^{d}$ (as long as $u_{\text{motion}}\cdot e_{d}=0$). To simplify the notation, we use a constant $u_{\text{motion}}$. Similarly the boundary condition could also be time-dependent, $u_{\varepsilon,BC}\in L^{\infty}(S;H^{1/2}(\Omega^{f}_{\varepsilon})^{d}\cap L^{\infty}(\Omega^{f}_{\varepsilon}))$, as this would not influence the following results.
@@
    The specific assumptions regarding $u_{BC}$ and, as a consequence, $u_{\varepsilon,BC}$ are a bit technical, but they ensure that there is an inflow and an outflow region and that $\Sigma$ and $\Gamma_{\varepsilon}$ are in relative motion to each other. The regularity assumption ensures compatibility across the different boundary parts.
@@
    The relationship between temperature and the viscosity of fluids is complex and often modeled heuristically. A commonly used example is the Vogel-Fulcher-Tammann equation, given by
@@
    where $\mu_{0},\alpha,T_{0}$ are empirical fitting parameters [5]. The model suggests that the relationship between viscosity and temperature tends to be smooth which means that the Lipschitz continuity is a reasonable assumption. However, this function $\mu$ is not bounded from above and therefore does not satisfy Assumption (A2). In practice, this law is only considered valid within a specific temperature range—namely, the interval where the material remains in its liquid state. Therefore, it makes sense to restrict this law to such a temperature range and extend it to $\mathbb{R}$ by its boundary values, which also preserves Lipschitz regularity. Moreover, the $L^{\infty}$-estimates we establish for the temperature field (see Lemma 6) provide a posteriori a qualitative justification for this restriction.
3 Analysis of the microscale model
@@ 3.1 Auxiliary results
    a linear extension operator $\mathbb{E}_{\varepsilon}^{f}:W^{1,p}(\Omega^{f}_{\varepsilon})\to W^{1,p}(\Omega_{\varepsilon})$ such that
    $\|\mathbb{E}_{\varepsilon}^{f}\psi\|_{W^{1,p}(\Omega_{\varepsilon})}\leq C\|\psi\|_{W^{1,p}(\Omega^{f}_{\varepsilon})}\quad\text{ for all }\psi\in W^{1,p}(\Omega^{f}_{\varepsilon}).$
    Here, $C>0$ may depend on $p$ but is independent of $\varepsilon$.
@@
    a linear extension operator $\mathbb{E}_{\varepsilon}^{s}:W^{1,p}(\Omega^{s}_{\varepsilon})\to W^{1,p}(\Omega)$ such that
    $\|\mathbb{E}_{\varepsilon}^{s}\varphi\|_{W^{1,p}(\Omega)}\leq C\|\varphi\|_{W^{1,p}(\Omega^{s}_{\varepsilon})}\quad\text{ for all }\varphi\in W^{1,p}(\Omega^{s}_{\varepsilon}),$
    with $C>0$ depended on $p$ but independent of $\varepsilon$.
@@
Both extension operators can be constructed using a classical extension procedure [43]. For thin layers, this was carried out in detail and further studied in [44, Theorem 1]. See also [36, Lemma A.6] for a generalization to other boundary conditions in the case $p=2$. The general idea is to first construct an extension operator for a single cell, in the case of the fluid from $W^{1,p}(Z)$ to $W^{1,p}(Y^{d})$. Which exists since $\Gamma$ is Lipschitz. The extension is $\varepsilon$-independent (as there is no $\varepsilon$ involved yet). The statement is then obtained by gluing these extension operators together via a periodic partition of unity in conjunction with a scaling argument. ∎
@@
    $\|\psi\|_{L^{p}(\Gamma_{\varepsilon})}^{p}\leq\frac{C}{\varepsilon}\left(\|\psi\|_{L^{p}(\Omega^{f}_{\varepsilon})}^{p}+\varepsilon^{p}\|\nabla\psi\|_{L^{p}(\Omega^{f}_{\varepsilon})}^{p}\right)\quad\text{for all}\quad\psi\in W^{1,p}(\Omega^{f}_{\varepsilon}).$
@@
    $\|\phi\|_{L^{p}(\Gamma_{\varepsilon})}\leq C\|\phi\|_{W^{1,p}(\Omega^{s}_{\varepsilon})}\quad\text{for all}\quad\varphi\in W^{1,p}(\Omega^{s}_{\varepsilon}).$
@@
$(i)$. This estimate for $\psi\in W^{1,p}(\Omega^{f}_{\varepsilon})$ is quite standard in homogenization and follows by a splitting argument and a change of variables, see for example [45, Lemma 5] for a more detailed explanation.
$(ii)$. This was proven in [36, Lemma A.5].
$(iii).$ The estimate can be derived by a scaling argument. With the extension operator from Lemma 1, it holds
$\|\psi\|_{L^{p}(\Sigma^{f}_{\varepsilon})}$ $=\|\mathbb{E}_{\varepsilon}^{f}\psi\|_{L^{p}(\Sigma^{f}_{\varepsilon})}\leq\|\mathbb{E}_{\varepsilon}^{f}\psi\|_{L^{p}(\partial\Sigma\times(0,\varepsilon))}=\varepsilon^{\frac{1}{p}}\|\mathbb{E}_{\varepsilon}^{f}\psi(\tilde{x},\varepsilon x_{d})\|_{L^{p}(\partial\Sigma\times(0,1))}$
$\leq C\varepsilon^{\frac{1}{p}}\|\mathbb{E}_{\varepsilon}^{f}\psi(\tilde{x},\varepsilon x_{d})\|_{W^{1,p}(\Omega)}\leq C\|\mathbb{E}_{\varepsilon}^{f}\psi\|_{W^{1,p}(\Omega_{\varepsilon})}\leq C\|\psi\|_{W^{1,p}(\Omega^{f}_{\varepsilon})}.$
Where we used the trace theorem on $\Omega$. Note, as before, the estimate could be improved by a $\varepsilon$ in front of the term with the $\partial_{x_{d}}$ norm, but this is not needed in the following analysis. ∎
@@
For any $\psi\in L^{\infty}(S;L^{2}(\Omega_{\varepsilon})\cap L^{2}(S;H^{1}(\Omega_{\varepsilon}))$, let $\bar{\psi}\in L^{\infty}(S;L^{2}(\Omega))\cap L^{2}(S;H^{1}(\Omega))$ be the corresponding upscaled function given by $\bar{\psi}(t,\tilde{x},x_{d})=\psi(t,\tilde{x},\varepsilon x_{d})$. Using standard interpolation embeddings (see, e.g., [37, Chapter 2, Inequality 3.8]), it holds